OpenStudy (anonymous):

The measure of angle 1 is five less than four times the measure of angle 2. If angle 1 and angle 2 form a linear pair, what are their measures?

OpenStudy (anonymous):

what'ss linear pair?

OpenStudy (anonymous):

Linear pair is also another expression that says that angle 1 and angle 2 are supplementary i.e. their measures add up to 180 degrees

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

so do I do 2(4x-5)=180?

OpenStudy (anonymous):

angle 1 = 4( angle 2) - 5 4( angle 2) - 5 + angle 2 = 180

OpenStudy (anonymous):

You need to write up two equation relating angle 1 and angle 2, one that is just saying that Angle 1 is equal to 4 times angle 2 minus 5, and another saying that they are supplementary, and solve the resulting linear system (by substitution or elimination)

OpenStudy (anonymous):

x = angle 1 y= angle 2 x = 4y - 5 x + y = 180 like that?

OpenStudy (anonymous):

right

OpenStudy (anonymous):